OpenStudy (anonymous):

what is the length of CD? A)5 B) 4 C)3 D) 6 HELP!!

OpenStudy (raden):

do you have a diagram ?

OpenStudy (anonymous):

here

OpenStudy (raden):

both triangles are similar. x/(20-x) = 7/21 solve for x CD is x

OpenStudy (anonymous):

the answer choices are A) 5 B) 4 c)3 D) 6

OpenStudy (raden):

x/(20-x) = 7/21 x/(20-x) = 1/3 cross multipy, get 3x = 20 -x .......... .......... .......... x = ... ?

OpenStudy (anonymous):

5!

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

can you help me on one more?please!

OpenStudy (raden):

3x = 20 -x 3x+x = 20 4x = 20 x = 20/4 = 5 you are right :)

OpenStudy (raden):

sure, post your question

OpenStudy (anonymous):

wwhat is the length of AC? A) 136 B) 72 C) 132 D) 96 !!!

OpenStudy (raden):

that is same like before. x/(144-x) = 3/51 solve for x, first then calculating the value of AC with AC = 144-x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dont understand :(!

OpenStudy (raden):

x/(144-x) = 3/51 x/(144-x) = 1/17 cross multiply, get 17x = 144-x ...... ...... x = ... ?

OpenStudy (raden):

what should you do, to get x if you have an equation : 17x = 144-x

OpenStudy (anonymous):

i got 9 but that doesnt match to my answer choices !

OpenStudy (raden):

well, see how to solving for x 17x = 144-x add by x to both sides, get 17x + x = 144-x+x 18x = 144 divided by 18 on both sides, get 18x/18 = 144/18 x = 8

OpenStudy (raden):

now, the question asking about AC. AC = 144 - x put x = 8, get AC = 144 - 8 = 136

OpenStudy (raden):

hoped that make sense

OpenStudy (anonymous):

thank u

OpenStudy (raden):

you're welcome